OpenStudy (anonymous):

find two numbers whose difference is -30 abd whose product is a minimum

OpenStudy (thejax):

Let x and y be the numbers. x-y=-30 Product = xy = (-30+y)-y

OpenStudy (debbieg):

x-y=-30 so x=-30+y Product = xy = (-30+y)*y which is a quadratic. Quadratic equations always have their minimum or maximum at the vertex (notice that this one is an upward opening parabola, so has a minimum). Find the vertex, and you have your y that minimizes the product; substitute that back into x=-30+y

OpenStudy (anonymous):

@DebbieG can you continue it, i can understand it little, sorry if i ask you so

OpenStudy (debbieg):

Well, I think I would have opted to solve the difference equation for y, instead of for x.... just so that you end up with a quadratic equation with y as a function of x, which is how we are typically used to seeing it. so let's start over. If the two numbers are x and y, you have: x-y=-30 so y=x + 30 (do you understand that step?) Now what we want to minimize is the product, xy. So substituting in the expression we obtained for y above, we have: Product = \(\large xy = x(x + 30) = x^2 + 30x\) So now the job is to find the x-value that minimizes \(\large f(x) = x^2 + 30x\) Since that is a quadratic function, it's graph is an upward opening parabola. So find the vertex, and you have found the minimum value of the function. Thus the x-coordinate of the vertex is where the function is at a minimum. Use the formula for the x-coordinate: \(\large x=\dfrac{-b}{2a}\) to find that x-coordinate. Once you have x, that is one of the numbers. Use the fact that \(\large y=x + 30\) to find y. And there are the two numbers.